OpenStudy (anonymous):

1 and 2step equations ...p-47=22 ? 9+s=-5 ? -2x=-8 ? y-(26)=41 ?

OpenStudy (saifoo.khan):

?

OpenStudy (anonymous):

?

OpenStudy (saifoo.khan):

lol, i cant get the problem

OpenStudy (anonymous):

wat do u mean cant get the poblem ?

OpenStudy (anonymous):

lol ??

OpenStudy (saifoo.khan):

i cant understand this.

OpenStudy (anonymous):

wat doing 1 step equations ???

OpenStudy (saifoo.khan):

Yeah.

OpenStudy (anonymous):

ooo lol i learned this but idont rememer how to do it

OpenStudy (saifoo.khan):

can u please rewrite the question?

OpenStudy (anonymous):

yeah np.

OpenStudy (saifoo.khan):

:D

OpenStudy (anonymous):

p-47=22

OpenStudy (anonymous):

9+s=-5

OpenStudy (anonymous):

-2-x=-8

OpenStudy (saifoo.khan):

p = 69 s = -14

OpenStudy (saifoo.khan):

x = 6

OpenStudy (saifoo.khan):

i cant solve it now.

OpenStudy (anonymous):

how did u figure thout ???

OpenStudy (saifoo.khan):

umm, idk.. they r the basic problems!

OpenStudy (anonymous):

oo lol bc i have to show my work ...

OpenStudy (anonymous):

ooooo so i add 47+22

OpenStudy (saifoo.khan):

Yeap!

OpenStudy (anonymous):

oooo ok :)

OpenStudy (anonymous):

y-(-26)=41

OpenStudy (saifoo.khan):

y + 26 = 41 y = 41 - 26 y = 15

OpenStudy (anonymous):

w/=3

OpenStudy (saifoo.khan):

what is taht?

OpenStudy (saifoo.khan):

that*

OpenStudy (anonymous):

ooo woopss i ment ... w/5 =3

OpenStudy (saifoo.khan):

w = 15

OpenStudy (saifoo.khan):

simply cross multiply.

OpenStudy (anonymous):

-24=8x

OpenStudy (saifoo.khan):

post them on LEFT>

OpenStudy (saifoo.khan):

x = -3

OpenStudy (anonymous):

o but how cud i writei to show my work

OpenStudy (saifoo.khan):

-24/3 = x

OpenStudy (saifoo.khan):

x = -3